For Toucan industries, the following relationships exists: each unit of output is sold for $35, the fixed costs are $160,000 and variable cost are $15 per unit.

A)  What is the firm’s gain or loss at sales of 6000 units and 9000 units?

 

B)   What is the break-even point? (Illustrate by means of a chart/graph)

 

 

C)   What is toucan’s degree of operating leverage at sales of 6000 units and 9000 units?

 

D)  What happens to the break-even point if the selling price rises to $40? What is the significance of the change to financial management? (Illustrate by means of a chart)

 

 

E)   What happens to the break-even point if the selling price rises to $40 but variable costs rise to $20 per unit?


1
Expert's answer
2021-06-17T13:04:35-0400

(a)At sales of 6000 units:

profit= total revenue - total cost

="(6000\times35)-(160000+(15\times6000))"

="-40000"

"\implies"loss of $40,000


At sales of 9000 units

profit= total revenue - total cost

"=(9000\times 35)-(160000+(15\times9000))"

"=20,000"

"\implies" profit of $20,000

(b) The break even point is the price point where marginal cost curve intersects average total cost curve.

"break even point=\frac{fixed costs}{total revenue-variable cost}"

="\frac {160,000}{315,000-135,000}"

"=0.89"



(c) Degree of operating leverages

at 6,000 units

"= \frac {sales-variable costs}{sales-variable costs -fixed costs}"


"= \frac {(6000\times35)-(6000\times15)}{(6000\times35)-(6000\times15)-160,000)}"


=-3


at 9000 units

"=\frac {(9000\times35)-(9000\times15)}{(9000\times35)-(9000\times15)-160,000}"


=9.

(d) if selling price rises to 40

"BEP= \frac{160,000}{(9000\times40)-135,000}"

= 0.71

BEP falls from 0.89 to 0.71

The decrease in BEP implies that there is a fall in the proportion of contribution margin products that are sold.

(e)

"BEP=\frac {160,000}{(9000\times40)-(9000\times20)}"

=0.89

break even point is raised back to the initial point 0.89.